3. Given TFC = $20, AVC =$10, and ATC = $11, find TVC and Q. (1.5 points)
Answer #1

Answer

AFC=ATC-AVC

=11-10

=1

Q=TFC/AFC=20/1=20

the Q is 20 units

TVC=AVC*Q

=10*20

=$200

the TVC is $200